Express the complex number in trigonometric form.

-3i

help me please
1
Expert's answer
2014-05-14T13:30:12-0400
Answer depends on the conventions of your textbook:
-3i=|z|(cos(alpha)+i*sin(alpha))=3(cos(-pi/2)+i*sin(-pi/2)) is expressed in trigonometric form when a polar angle ranges from -pi to pi;
-3i=|z|(cos(alpha)+i*sin(alpha))=3(cos(3*pi/2)+i*sin(3*pi/2)) is expressed in trigonometric form when a polar angle ranges from 0 to 2*pi.